OpenStudy (anonymous):

Find the sum of the x- and y-intercepts of any tangent line to the curve √x+√y=√7.

OpenStudy (anonymous):

i used implicit differentiation to get dy/dx= -√y/√x

OpenStudy (shubhamsrg):

yep,,now let x1,y1 be the point on the curve, eqn of tangent at that point is then y-y1 = -sqrt(y1/x1) (x-x1) convert it into the form of y/a + x/b =1 a and b are y and x intercepts respectively.. just substitute in a+b

OpenStudy (anonymous):

ahhh so the answer is 7! thanks got it

OpenStudy (shubhamsrg):

7 ?

OpenStudy (shubhamsrg):

ohh..maybe,,alright.. you'll also have to use the fact sqrt(x1) + sqrt(y1) = sqrt(7) 7 might be the ans ,,i dont know..

OpenStudy (anonymous):

ya it is thanks